OpenStudy (anonymous):

A rectangular prism has a length of 12 cm and a width of 4 cm. If the volume of the prism is 240 cm3, what is the height?

OpenStudy (anonymous):

240/(12 * 4) = 5cm

OpenStudy (anonymous):

Thats wrong, the volume is got by first finding the area of the triangle and then multiplying by the width.

OpenStudy (anonymous):

The volume of a rectangular prism can be found by the formula: volume=length*width*height 240 = 12 * 4 * height height = 240 /(12 * 4) = 5cm FORGIVE ME FOR SKIPPING STEPS

OpenStudy (anonymous):

http://math.about.com/od/formulas/ss/surfaceareavol_4.htm Are nor triangles I agree with first solution

OpenStudy (anonymous):

No the answer is still wrong .. the area is 12*4 /2 then multiply by height! I hope you get the point!

OpenStudy (anonymous):

where are you getting a triangle? the problem is stated as a rectangular prism

OpenStudy (anonymous):

SORRRRYYY MY BAD!!! REALLY SORRY! Its correct